OpenStudy (anonymous):

What is the lateral area of a rectangular prism if the base edges are 8 feet and 6 feet and the height is 10 feet? Explain lateral area to me

Parth (parthkohli):

\( \color{Black}{\Rightarrow la = 2h(w +l)}\)

Parth (parthkohli):

\( \color{Black}{\Rightarrow la = 2(10) \times (8 + 6)}\)

Parth (parthkohli):

\( \color{Black}{\Rightarrow la = 20 \times 14}\)

OpenStudy (anonymous):

Thank you alot

Parth (parthkohli):

You're welcome!

OpenStudy (compassionate):

 (a, b, and c are the lengths of the 3 sides) In words, the surface area of a rectangular prism is the area of the six rectangles that cover it. But we don't have to figure out all six because we know that the top and bottom are the same, the front and back are the same, and the left and right sides are the same. The area of the top and bottom (side lengths a and c) = a*c. Since there are two of them, you get 2ac. The front and back have side lengths of b and c. The area of one of them is b*c, and there are two of them, so the surface area of those two is 2bc. The left and right side have side lengths of a and b, so the surface area of one of them is a*b. Again, there are two of them, so their combined surface area is 2ab.

Parth (parthkohli):

@Compassionate http://www.math.com/tables/geometry/surfareas.htm#rec You just copied it from here.

OpenStudy (compassionate):

Yes, I did. I should have cited. I thought that this website would give him a good idea of lateral surface area. There is nothing wrong with what I did, please do not call me out as if I did something wrong.

Parth (parthkohli):

Ah! No, you didn't. You had good intentions.

OpenStudy (compassionate):

Oh, I'm sorry for the misunderstanding, Parth. I'm sorry. I'll make sure to cite next time, though.

OpenStudy (anonymous):

is that form of the formula used for all shapes?

Parth (parthkohli):

Nah, it's all cool.

Parth (parthkohli):

Nope, Mastercat.

Parth (parthkohli):

The lateral area of a cylinder is \(2\pi r \times h\).

OpenStudy (compassionate):

No, Mastercat. Each shape has a different formula. I suggest you Google them.

OpenStudy (anonymous):

damn theres always a catch..

Parth (parthkohli):

Exactly — and that is why I don't like Geometry.

OpenStudy (anonymous):

So what is the form for triangles?

Parth (parthkohli):

Triangles are not three-dimensional, so they don't have a lateral area. Still - the area of a triangle is \(\text{base } \times \text{altitude}\).

OpenStudy (anonymous):

I mean pyramids*

OpenStudy (compassionate):

Mastercat, see this chart for all the surface area formulas. http://www.mathwords.com/s/surface_area.htm

OpenStudy (anonymous):

<3 #lifesaved

Parth (parthkohli):

Hmm, the lateral area is different from surface area.

OpenStudy (anonymous):

yeah i need the lateral area ones now.

OpenStudy (anonymous):

I guess I shall look.

Parth (parthkohli):

Sure.